OpenStudy (anonymous):

NEED SOME ONE TO CHECK MY ANSWER 1/3^-5 =243

OpenStudy (anonymous):

@Nnesha

Nnesha (nnesha):

why do you thin it's 243 instead 1/243 ?? :=)

OpenStudy (anonymous):

MY KEYBORD IS BROK

Nnesha (nnesha):

okay.so is that the answer of my question ?? :=) btw buy new one :PD

OpenStudy (anonymous):

LOL

OpenStudy (anonymous):

R U SURE THAT IS HOW I WRITE IT

Nnesha (nnesha):

oky. so my question is why it is not 1/243 ?

OpenStudy (anonymous):

IDK THAT IS THE ANSWER I CAME UP WIT

Nnesha (nnesha):

Can i ask you a question ?

OpenStudy (anonymous):

YEA

Nnesha (nnesha):

how did you get 243 ?

OpenStudy (anonymous):

1÷(3^(-5))=243

OpenStudy (anonymous):

YEA

OpenStudy (texaschic101):

is the question (1/3)^-5 or is it 1/(3^-5)

Nnesha (nnesha):

if you see negative exponents just `flip` the fraction \[\huge\rm x^{-m}=\frac{ 1 }{ x^m }\]

Nnesha (nnesha):

\[\huge\rm \frac{ x^{-m} }{ 1 }=\frac{ 1 }{ x^m }\]

OpenStudy (anonymous):

THAT HOW IT LOOKS ON MY SCREEN

Nnesha (nnesha):

http://prntscr.com/893jr7 buy this one :=)

OpenStudy (texaschic101):

lol...apple computer picture...thats a good one NNesha

OpenStudy (anonymous):

LOL

Nnesha (nnesha):

thank you thank you! ;P

OpenStudy (texaschic101):

(1/3) ^ -5 = 1/(1/3^5) = 1/(1/3*1/3*1/3*1/3*1/3) = 1/243

OpenStudy (anonymous):

OK

Nnesha (nnesha):

i guess it's \[\frac{ 1 }{ 3^{-5} }\]